THE MOMENTUM OF A PARTICLE MOVING IN STRAIGHT LINE IS GIVEN BY P =LN T+ I BY T KGM BY S THE TIMET IS   grater than 0 at which the net force acting on particle is 0 and its momentum at that time is

Grade:9

2 Answers

Arun
4 years ago

F=dp/dt

therefore F=1/t - 1/t^2

F=0........t=1

momentum=1

 

In case of any difficulty, please feel free to ask again.

 

Regards

Arun (askIITians forum expert)

Khimraj
4 years ago
p = lnt + 1/t
F = dp/dt = 1/t – 1/t2
F = 0 = t2 -t
t=1
p = ln1 + 1 =1
…........................................................................................